# Dirichlet, Neumann, Mixed and self-dual holography:

### v1

@@ 3.3 AdS/CFT dictionary
Lastly, as it was mentioned, Dirichlet and Neumann results are closely related, which can be seen at the level of bulk-to-bulk $G$ and boundary-to-bulk $K$ propagators. With the help of the bulk-to-bulk propagators given in Appendix B one can verify the following relations [^12] Here, $\Pi^{\text{e}}$ and $\Pi^{\text{o}}$ are projectors onto the parity-even and parity-odd structures of the current’s two-point functions, see Appendix A.
@@ 4 Correlators
Essential ingredients are bulk-to-bulk propagators, which we collect in Appendix B. [^13] Propagators with mixed boundary conditions in Feynman and axial gauges in YM and Chalmers–Siegel theory seem new, from which SDYM’s propagators follow. In the gauge $A_{3,0}=0$ the spin-one propagator for mixed boundary conditions was found in [43]. External lines in Feynman-Witten diagrams are saturated by boundary-to-bulk propagators, which can be derived from the bulk-to-bulk ones. The kinematics of the boundary-to-bulk propagators is fixed and is visible already in the Fefferman-Graham expansion of Section 3.1. However, there are simple prefactors that depend on boundary conditions. Concretely, $\Phi$ and $\Psi$ external lines in Witten diagrams are to be replaced by
@@ 4.1 Two-point functions
With the help of Appendix B, one finds
@@ 4.3.2 SDYM Axial
with $\Delta\eta^{\text{L}}_{AA}$ defined in (B.14). It is only this difference that we compute, i.e.
@@ 5 Discussion and Conclusions
At least when a theory is conformally invariant, e.g. SDYM and higher spin extensions of it, there can be an interesting correspondence between celestial and AdS/CFT structures, see e.g. [62]. It would also be important to reconsider the vanishing of the leading piece of the AdS/CFT correlators in self-dual theories due to the recent discovery [63] that (++)(+...+-)-amplitudes are supported on the collinear kinematics. This observation seems to have a higher-spin extension [64, 65], where chiral higher-spin gravity in flat space is dual to the flat space version of the free vector model, which is a celestial analog of the higher-spin/vector model AdS/CFT duality [66, 67, 30], one major difference being that both sides of the duality are chiral.
Coming back to SDYM, in order to take $i\epsilon$’s in the propagators properly into account one has to return to Minkowski signature. However, at the level of three-point functions, the net effect is just a change of variables. Indeed, let us take (4.16) and set $q_{1}=q_{2}=r$
$-\frac{2g}{E}\frac{\langle q_{1}3\rangle\langle q_{2}3\rangle}{\langle q_{1}1\rangle\langle q_{2}2\rangle}\langle\bar{1}\bar{2}\rangle$ $=-\frac{2g}{E}\frac{\langle r3\rangle^{2}}{\langle r1\rangle\langle r2\rangle}\langle\bar{1}\bar{2}\rangle\,.$ (5.1)
$\delta^{4}(\sum_{i}i_{A}\bar{i}_{A^{\prime}})$ $=\langle\mu\nu\rangle^{2}\delta^{2}(\sum_{i}\langle\mu i\rangle\bar{i}_{A^{\prime}})\delta^{2}(\sum_{i}\langle\nu i\rangle\bar{i}_{A^{\prime}})\,.$ (5.2)
$\delta^{2}(\sum_{i}\langle\nu i\rangle\bar{i}_{A^{\prime}})$ $=|\langle\bar{1}\bar{2}\rangle|^{-1}\delta(\langle 13\rangle)\delta(\langle 23\rangle)\,.$ (5.3)
$\frac{\langle r3\rangle^{4}}{\langle r1\rangle\langle r2\rangle}\operatorname{sign}\langle\bar{1}\bar{2}\rangle\,\delta(\langle 13\rangle)\,\delta(\langle 23\rangle)\,\delta^{2}(\sum_{i}\langle ri\rangle\bar{i}_{A^{\prime}})\,.$ (5.4)
$\langle r3\rangle^{2}\frac{\operatorname{sign}\langle\bar{1}\bar{2}\rangle}{\langle\bar{1}\bar{2}\rangle}\,\delta\left(\langle 13\rangle-\tfrac{E\langle\bar{2}3\rangle}{\langle\bar{1}\bar{2}\rangle}\right)\,\delta\left(\langle 23\rangle-\tfrac{E\langle 3\bar{1}\rangle}{\langle\bar{1}\bar{2}\rangle}\right)\,\delta^{2}\left(\sum_{i}\langle ri\rangle\bar{i}_{A^{\prime}}-r_{A^{\prime}}E\right)\,$, (5.5)
which can now be combined with (4.18).
@@ Appendix A Spinor helicity and CFT
Appendix B Propagators
@@
$G(-k,z;k,z^{\prime})$ $=G^{\text{hom}}(-k,z;k,z^{\prime})+G^{\text{inh}}(-k,z;k,z^{\prime})\,.$ (B.1)
@@
$\begin{aligned} e^{i\gamma}\langle\Psi^{AA}(-k,z\rightarrow 0)\Psi^{BB}(k,z^{\prime})\rangle&=-e^{-i\gamma}\langle\bar{\Psi}^{AA}(-k,z\rightarrow 0)\Psi^{BB}(k,z^{\prime})\rangle\,,\\ e^{i\gamma}\langle\Psi^{AA}(-k,z\rightarrow 0)\bar{\Psi}^{BB}(k,z^{\prime})\rangle&=-e^{-i\gamma}\langle\bar{\Psi}^{AA}(-k,z\rightarrow 0)\bar{\Psi}^{BB}(k,z^{\prime})\rangle\,.\end{aligned}$ (B.3)
@@
$e^{i\gamma}\langle\Psi^{AA}(-k,z\rightarrow 0)\Phi^{B,B^{\prime}}(k,z^{\prime})\rangle$ $=-e^{-i\gamma}\langle\bar{\Psi}^{AA}(-k,z\rightarrow 0)\Phi^{B,B}(k,z^{\prime})\rangle\,.$ (B.4)
@@ B.1 YM: Feynman
$\langle\Phi_{A,A^{\prime}}(-k,z)\Phi_{B,B^{\prime}}(k,z^{\prime})\rangle_{\text{D/N}}=-\frac{\epsilon_{AB}\epsilon_{A^{\prime}B^{\prime}}}{2k}e^{-k|z-z^{\prime}|}\pm e^{-k(z+z^{\prime})}\frac{1}{2k}\left(\epsilon_{AB}\epsilon_{A^{\prime}B^{\prime}}-\epsilon_{AA^{\prime}}\epsilon_{BB^{\prime}}\right)\,$, (B.5)
@@
$\begin{aligned} \langle\Phi_{A,A^{\prime}}(-k,z)\Phi_{B,B^{\prime}}(k,z^{\prime})\rangle_{\gamma}&=\frac{\cos\gamma}{k}\Big(\cos{\gamma}\,\Pi^{\text{e}}_{AA^{\prime},BB^{\prime}}+i\sin{\gamma}\,\Pi^{\text{o}}_{AA^{\prime},BB^{\prime}}\Big)e^{-k(z+z^{\prime})}\,.\end{aligned}$ (B.7)
Note that in Feynman gauge, the Dirichlet propagator is not part of the continuous family of mixed boundary conditions, as setting $\gamma=0$ does not yield the Dirichlet propagator. Instead, one has to pick the Dirichlet boundary condition by hand from (B.5). Meanwhile, the Neumann propagator is smoothly obtained by setting $\gamma=\pi/2$.
B.2 YM: Axial
The propagator in axial gauge can be related to the Feynman gauge propagator through (B.2). We denote the difference between the Feynman gauge and axial gauge propagator as $\Delta\langle\Phi\Phi\rangle$. We note that the $\gamma$-dependent component of the propagator is the same for Feynman and axial gauge, so it suffices to look at the gauge variation of the Dirichlet and Neumann propagators. This is given by
@@ B.2 YM: Axial
B.3 Chalmers–Siegel: Feynman
@@
$\langle\Psi_{AA}(-k,z)\Phi_{B,B^{\prime}}(k,z^{\prime})\rangle=\nabla^{-k,z}_{AA^{\prime}}\langle\Phi{}_{A,}{}^{A^{\prime}}\,(-k,z)\Phi_{B,B^{\prime}}(k,z^{\prime})\rangle.$ (B.9)
@@
$\begin{aligned} \langle\Psi_{AA}(-k,z)\Phi_{B,B^{\prime}}(k,z^{\prime})\rangle_{\text{D}}&=-\frac{\epsilon_{AB}}{2k}\big(k_{AB^{\prime}}-\text{sign}(z-z^{\prime})k\epsilon_{AB^{\prime}}\big)e^{-k|z-z^{\prime}|}+\frac{\epsilon_{AB^{\prime}}}{2k}\bar{k}_{A}k_{B}e^{-k(z+z^{\prime})}\,,\\ \langle\Psi_{AA}(-k,z)\Phi_{B,B^{\prime}}(k,z^{\prime})\rangle_{\text{N}}&=-\frac{\epsilon_{AB}}{2k}\big(k_{AB^{\prime}}-\text{sign}(z-z^{\prime})k\epsilon_{AB^{\prime}}\big)e^{-k|z-z^{\prime}|}-\frac{\epsilon_{AB^{\prime}}}{2k}\bar{k}_{A}k_{B}e^{-k(z+z^{\prime})}\,,\\ \langle\Psi_{AA}(-k,z)\Phi_{B,B^{\prime}}(k,z^{\prime})\rangle_{\gamma}&=-\frac{1+e^{-2i\gamma}}{4k^{2}}\bar{k}_{A}\bar{k}_{A}k_{B}k_{B^{\prime}}e^{-k(z+z^{\prime})}\,.\end{aligned}$ (B.10)
@@
It consists of a term that satisfies (B.11) and another term containing a delta-function, which does not satisfy the equation of motion. The actual $\langle\Psi\Psi\rangle$ propagators are
$\begin{aligned} \langle\Psi_{AA}(-k,z)\Psi_{BB}(k,z^{\prime})\rangle_{\text{D}}&=\frac{\bar{k}_{A}\bar{k}_{A}k_{B}k_{B}}{2k}e^{-k(z+z^{\prime})}\,,\\ \langle\Psi_{AA}(-k,z)\Psi_{BB}(k,z^{\prime})\rangle_{\text{N}}&=-\frac{\bar{k}_{A}\bar{k}_{A}k_{B}k_{B}}{2k}e^{-k(z+z^{\prime})}\,,\\ \langle\Psi_{AA}(-k,z)\Psi_{BB}(k,z^{\prime})\rangle_{\gamma}&=\frac{1+e^{-2i\gamma}}{2k}\bar{k}_{A}\bar{k}_{A}k_{B}k_{B}e^{-k(z+z^{\prime})}\,.\end{aligned}$ (B.13)
B.4 Chalmers–Siegel: Axial
@@
where $\eta^{\text{L}}_{AA}\equiv\nabla^{-k,z}_{AA^{\prime}}\xi^{\text{L}}{}_{A}{}^{A^{\prime}}\,$ and $\xi^{\text{L}}_{AA^{\prime}}$ is defined in (B.8).
B.5 SDYM: Feynman

### v2

@@ 3.3 AdS/CFT dictionary
Lastly, as it was mentioned, Dirichlet and Neumann results are closely related, which can be seen at the level of bulk-to-bulk $G$ and boundary-to-bulk $K$ propagators. With the help of the bulk-to-bulk propagators given in Appendix C one can verify the following relations [^12] Here, $\Pi^{\text{e}}$ and $\Pi^{\text{o}}$ are projectors onto the parity-even and parity-odd structures of the current’s two-point functions, see Appendix A.
@@ 4 Correlators
Essential ingredients are bulk-to-bulk propagators, which we collect in Appendix C. [^13] Propagators with mixed boundary conditions in Feynman and axial gauges in YM and Chalmers–Siegel theory seem new, from which SDYM’s propagators follow. In the gauge $A_{3,0}=0$ the spin-one propagator for mixed boundary conditions was found in [43]. External lines in Feynman-Witten diagrams are saturated by boundary-to-bulk propagators, which can be derived from the bulk-to-bulk ones. The kinematics of the boundary-to-bulk propagators is fixed and is visible already in the Fefferman-Graham expansion of Section 3.1. However, there are simple prefactors that depend on boundary conditions. Concretely, $\Phi$ and $\Psi$ external lines in Witten diagrams are to be replaced by
@@ 4.1 Two-point functions
With the help of Appendix C, one finds
@@ 4.3.2 SDYM Axial
with $\Delta\eta^{\text{L}}_{AA}$ defined in (C.14). It is only this difference that we compute, i.e.
@@ 5 Discussion and Conclusions
At least when a theory is conformally invariant, e.g. SDYM and higher spin extensions of it, there can be an interesting correspondence between celestial and AdS/CFT structures, see e.g. [62]. It would also be important to reconsider the vanishing of the leading piece of the AdS/CFT correlators in self-dual theories due to the recent discovery [63] that (++)(+...+-)-amplitudes are supported on the collinear kinematics. This observation seems to have a higher-spin extension [64, 65], where chiral higher-spin gravity in flat space is dual to the flat space version of the free vector model, which is a celestial analog of the higher-spin/vector model AdS/CFT duality [66, 67, 30], one major difference being that both sides of the duality are chiral. We make a number of comments on the collinear limit of AdS/CFT correlators in Appendix B.
@@ Appendix A Spinor helicity and CFT
Appendix B Collinear limit
In order to take $i\epsilon$’s in the propagators properly into account one has to return to Minkowski signature. However, at the level of three-point functions, the net effect is just a change of variables. Indeed, let us take (4.16) and set $q_{1}=q_{2}=r$
$-\frac{2g}{E}\frac{\langle q_{1}3\rangle\langle q_{2}3\rangle}{\langle q_{1}1\rangle\langle q_{2}2\rangle}\langle\bar{1}\bar{2}\rangle$ $=-\frac{2g}{E}\frac{\langle r3\rangle^{2}}{\langle r1\rangle\langle r2\rangle}\langle\bar{1}\bar{2}\rangle\,.$ (B.1)
$\delta^{4}(\sum_{i}i_{A}\bar{i}_{A^{\prime}})$ $=\langle\mu\nu\rangle^{2}\delta^{2}(\sum_{i}\langle\mu i\rangle\bar{i}_{A^{\prime}})\delta^{2}(\sum_{i}\langle\nu i\rangle\bar{i}_{A^{\prime}})\,.$ (B.2)
$\delta^{2}(\sum_{i}\langle\nu i\rangle\bar{i}_{A^{\prime}})$ $=|\langle\bar{1}\bar{2}\rangle|^{-1}\delta(\langle 13\rangle)\delta(\langle 23\rangle)\,.$ (B.3)
$\frac{\langle r3\rangle^{4}}{\langle r1\rangle\langle r2\rangle}\operatorname{sign}\langle\bar{1}\bar{2}\rangle\,\delta(\langle 13\rangle)\,\delta(\langle 23\rangle)\,\delta^{2}(\sum_{i}\langle ri\rangle\bar{i}_{A^{\prime}})\,.$ (B.4)
$\langle r3\rangle^{2}\frac{\operatorname{sign}\langle\bar{1}\bar{2}\rangle}{\langle\bar{1}\bar{2}\rangle}\,\delta\left(\langle 13\rangle-\tfrac{E\langle\bar{2}3\rangle}{\langle\bar{1}\bar{2}\rangle}\right)\,\delta\left(\langle 23\rangle-\tfrac{E\langle 3\bar{1}\rangle}{\langle\bar{1}\bar{2}\rangle}\right)\,\delta^{2}\left(\sum_{i}\langle ri\rangle\bar{i}_{A^{\prime}}-r_{A^{\prime}}E\right)\,$, (B.5)
which can now be combined with (4.18) to get the complete three-point function.
Starting from the five-point functions the collinear limit becomes highly nontrivial [63]. At the four-point level the collinear limit in flat space gets a simple contribution from the “cut” in accordance with $1/(p^{2}+i\epsilon)=\mathrm{PV}(1/p^{2})-i\pi\delta(p^{2})$. A careful analysis of the AdS-propagator is needed to recover the $i\epsilon$’s. A shortcut is to consider to the flat limit:
$\int e^{-(k_{1}+k_{2})z}\langle\Phi_{AA^{\prime}}(-k,z)\Phi_{BB^{\prime}}(k,z^{\prime})\rangle e^{-(k_{3}+k_{4})z^{\prime}}=\frac{1}{E}\big(-\frac{\epsilon_{AB}\epsilon_{A^{\prime}B^{\prime}}}{E_{\text{L}}E_{\text{R}}-i\epsilon}\big)+\dots\,.$ (B.6)
This leads to the following $\delta$-function,
$\delta(E_{\text{L}}E_{\text{R}})=\delta(\langle 12\rangle\langle\bar{1}\bar{2}\rangle-EE_{\text{L}})=\delta(\langle 34\rangle\langle\bar{3}\bar{4}\rangle-EE_{\text{R}})=\frac{1}{|\langle\bar{3}\bar{4}\rangle|}\delta(\langle 34\rangle-E\frac{E_{\text{R}}}{\langle\bar{3}\bar{4}\rangle})\,.$ (B.7)
The momentum conservation can be treated as before. We first insert $E$ to restore the four-momentum conservation and write
$\delta^{4}(\sum\limits_{i=0}^{4}i_{A}\bar{i}_{A^{\prime}})=\delta^{2}(\sum\limits_{i=0}^{4}\langle ri\rangle\bar{i}_{A^{\prime}}-r_{A^{\prime}}E)\delta^{2}(\sum\limits_{i=0}^{4}\langle 4i\rangle\bar{i}_{A^{\prime}}-4_{A^{\prime}}E)\langle r4\rangle^{2}\,.$ (B.8)
The second delta leads to
$\delta^{2}(\sum\limits_{i=0}^{4}\langle 4i\rangle\bar{i}_{A^{\prime}}-4_{A^{\prime}}E)=\delta\Big(\langle 14\rangle-\frac{E}{\langle\bar{1}\bar{2}\rangle}\big(E_{\text{R}}\frac{\langle\bar{2}\bar{3}\rangle}{\langle\bar{3}\bar{4}\rangle}+\langle\bar{2}4\rangle\big)\Big)\delta\Big(\langle 24\rangle+\frac{E}{\langle\bar{1}\bar{2}\rangle}\big(E_{\text{R}}\frac{\langle\bar{1}\bar{3}\rangle}{\langle\bar{3}\bar{4}\rangle}+\langle\bar{1}4\rangle\Big)\,.$ (B.9)
Now, everything can be combined with the four-point functions in the main text.
Appendix C Propagators
@@
$G(-k,z;k,z^{\prime})$ $=G^{\text{hom}}(-k,z;k,z^{\prime})+G^{\text{inh}}(-k,z;k,z^{\prime})\,.$ (C.1)
@@
$\begin{aligned} e^{i\gamma}\langle\Psi^{AA}(-k,z\rightarrow 0)\Psi^{BB}(k,z^{\prime})\rangle&=-e^{-i\gamma}\langle\bar{\Psi}^{AA}(-k,z\rightarrow 0)\Psi^{BB}(k,z^{\prime})\rangle\,,\\ e^{i\gamma}\langle\Psi^{AA}(-k,z\rightarrow 0)\bar{\Psi}^{BB}(k,z^{\prime})\rangle&=-e^{-i\gamma}\langle\bar{\Psi}^{AA}(-k,z\rightarrow 0)\bar{\Psi}^{BB}(k,z^{\prime})\rangle\,.\end{aligned}$ (C.3)
@@
$e^{i\gamma}\langle\Psi^{AA}(-k,z\rightarrow 0)\Phi^{B,B^{\prime}}(k,z^{\prime})\rangle$ $=-e^{-i\gamma}\langle\bar{\Psi}^{AA}(-k,z\rightarrow 0)\Phi^{B,B}(k,z^{\prime})\rangle\,.$ (C.4)
@@ C.1 YM: Feynman
$\langle\Phi_{A,A^{\prime}}(-k,z)\Phi_{B,B^{\prime}}(k,z^{\prime})\rangle_{\text{D/N}}=-\frac{\epsilon_{AB}\epsilon_{A^{\prime}B^{\prime}}}{2k}e^{-k|z-z^{\prime}|}\pm e^{-k(z+z^{\prime})}\frac{1}{2k}\left(\epsilon_{AB}\epsilon_{A^{\prime}B^{\prime}}-\epsilon_{AA^{\prime}}\epsilon_{BB^{\prime}}\right)\,$, (C.5)
@@
$\begin{aligned} \langle\Phi_{A,A^{\prime}}(-k,z)\Phi_{B,B^{\prime}}(k,z^{\prime})\rangle_{\gamma}&=\frac{\cos\gamma}{k}\Big(\cos{\gamma}\,\Pi^{\text{e}}_{AA^{\prime},BB^{\prime}}+i\sin{\gamma}\,\Pi^{\text{o}}_{AA^{\prime},BB^{\prime}}\Big)e^{-k(z+z^{\prime})}\,.\end{aligned}$ (C.7)
Note that in Feynman gauge, the Dirichlet propagator is not part of the continuous family of mixed boundary conditions, as setting $\gamma=0$ does not yield the Dirichlet propagator. Instead, one has to pick the Dirichlet boundary condition by hand from (C.5). Meanwhile, the Neumann propagator is smoothly obtained by setting $\gamma=\pi/2$.
C.2 YM: Axial
The propagator in axial gauge can be related to the Feynman gauge propagator through (C.2). We denote the difference between the Feynman gauge and axial gauge propagator as $\Delta\langle\Phi\Phi\rangle$. We note that the $\gamma$-dependent component of the propagator is the same for Feynman and axial gauge, so it suffices to look at the gauge variation of the Dirichlet and Neumann propagators. This is given by
@@ C.2 YM: Axial
C.3 Chalmers–Siegel: Feynman
@@
$\langle\Psi_{AA}(-k,z)\Phi_{B,B^{\prime}}(k,z^{\prime})\rangle=\nabla^{-k,z}_{AA^{\prime}}\langle\Phi{}_{A,}{}^{A^{\prime}}\,(-k,z)\Phi_{B,B^{\prime}}(k,z^{\prime})\rangle.$ (C.9)
@@
$\begin{aligned} \langle\Psi_{AA}(-k,z)\Phi_{B,B^{\prime}}(k,z^{\prime})\rangle_{\text{D}}&=-\frac{\epsilon_{AB}}{2k}\big(k_{AB^{\prime}}-\text{sign}(z-z^{\prime})k\epsilon_{AB^{\prime}}\big)e^{-k|z-z^{\prime}|}+\frac{\epsilon_{AB^{\prime}}}{2k}\bar{k}_{A}k_{B}e^{-k(z+z^{\prime})}\,,\\ \langle\Psi_{AA}(-k,z)\Phi_{B,B^{\prime}}(k,z^{\prime})\rangle_{\text{N}}&=-\frac{\epsilon_{AB}}{2k}\big(k_{AB^{\prime}}-\text{sign}(z-z^{\prime})k\epsilon_{AB^{\prime}}\big)e^{-k|z-z^{\prime}|}-\frac{\epsilon_{AB^{\prime}}}{2k}\bar{k}_{A}k_{B}e^{-k(z+z^{\prime})}\,,\\ \langle\Psi_{AA}(-k,z)\Phi_{B,B^{\prime}}(k,z^{\prime})\rangle_{\gamma}&=-\frac{1+e^{-2i\gamma}}{4k^{2}}\bar{k}_{A}\bar{k}_{A}k_{B}k_{B^{\prime}}e^{-k(z+z^{\prime})}\,.\end{aligned}$ (C.10)
@@
It consists of a term that satisfies (C.11) and another term containing a delta-function, which does not satisfy the equation of motion. The actual $\langle\Psi\Psi\rangle$ propagators are
$\begin{aligned} \langle\Psi_{AA}(-k,z)\Psi_{BB}(k,z^{\prime})\rangle_{\text{D}}&=\frac{\bar{k}_{A}\bar{k}_{A}k_{B}k_{B}}{2k}e^{-k(z+z^{\prime})}\,,\\ \langle\Psi_{AA}(-k,z)\Psi_{BB}(k,z^{\prime})\rangle_{\text{N}}&=-\frac{\bar{k}_{A}\bar{k}_{A}k_{B}k_{B}}{2k}e^{-k(z+z^{\prime})}\,,\\ \langle\Psi_{AA}(-k,z)\Psi_{BB}(k,z^{\prime})\rangle_{\gamma}&=\frac{1+e^{-2i\gamma}}{2k}\bar{k}_{A}\bar{k}_{A}k_{B}k_{B}e^{-k(z+z^{\prime})}\,.\end{aligned}$ (C.13)
C.4 Chalmers–Siegel: Axial
@@
where $\eta^{\text{L}}_{AA}\equiv\nabla^{-k,z}_{AA^{\prime}}\xi^{\text{L}}{}_{A}{}^{A^{\prime}}\,$ and $\xi^{\text{L}}_{AA^{\prime}}$ is defined in (C.8).
C.5 SDYM: Feynman